OpenStudy (anonymous):

how to you find the range of f(x)=x-4/x^2-4 PLEASE HELP

OpenStudy (anonymous):

\[\frac{x-4}{x^2-4}\]?

OpenStudy (anonymous):

calculus problem or something else?

OpenStudy (anonymous):

guess i'll never know...

OpenStudy (anonymous):

sorry! its pre cal and yest thats correct!

OpenStudy (anonymous):

ok without calc it is not going to be that easy

OpenStudy (anonymous):

do you have any insight on how to solve i t?

OpenStudy (anonymous):

yes, and it is a pain

OpenStudy (anonymous):

we are going to have to solve for x

OpenStudy (anonymous):

\[y=\frac{x-4}{x^2-4}\] \[yx^2-4y=x-4\] \[yx^2-x-4y-4=0\] now we have a quadratic equation in x, so we can solve using the quadratic formula

OpenStudy (mertsj):

\[x=\frac{1\pm \sqrt{1+16y^2+16y}}{2y}\]

OpenStudy (anonymous):

@Mertsj if you have a snap way to do it let me know

OpenStudy (anonymous):

ok we are doing the same thing

OpenStudy (mertsj):

Graph it.

OpenStudy (anonymous):

denominator is y^2 but it makes no difference now you have to solve \[16x^2+16x+1\geq 0\] lord!

OpenStudy (anonymous):

oh no you are right, denominator is 2y

OpenStudy (mertsj):

How so? a = y doesn't it?

OpenStudy (anonymous):

so we have to solve [16x^2+16y+1=0\] using the quadratic formula, another pain oh yes you are right, i am wrong about the denominator

OpenStudy (anonymous):

\[16x^2+16y+1=0\]

OpenStudy (anonymous):

\[x=\frac{-2\pm\sqrt{3}}{4}\]

OpenStudy (mertsj):

I got 8 on the bottom

OpenStudy (anonymous):

also wrong, sorry range is \[(-\infty,\frac{-2-\sqrt{3}}{4}]\cup (\frac{-2+\sqrt{3}}{4},\infty)\]

OpenStudy (anonymous):

well i have to admit i cheated for the last part http://www.wolframalpha.com/input/?i=16x^2%2B16x%2B1%3E0

OpenStudy (mertsj):

I know. I found it.

OpenStudy (mertsj):

Do you really think a pre-calc student was supposed to do that?